Parth (parthkohli):

Why is \(0 \times 0 \) not \(\text{undefined}\)? How can we multiply nothing to nothing?

Parth (parthkohli):

\(0 \times 1\), on the other hand, can be calculated by adding 0 to itself once, which leaves 0. Same for all other \(0 \times n\) where \(n\) is not zero.

Parth (parthkohli):

How can we add 0 to itself no number of times?

OpenStudy (anonymous):

0=1-1 0*0=0*(1-1)=0-0=0

OpenStudy (anonymous):

zero is an approximation anyway.

Parth (parthkohli):

There still can be mathematical fallacy in any proof, @apple-pi.